Not even sure where to begin this one. Any help would be appreciated.
Find an equation for the ellipse whose graph goes through the points (16,0) and (8,6). If your answer is (x^2)/A + (y^2)/B = 1, then A=_____ and B=_____
Thanks in advance for your help.

Answer by stanbon(75887)

Find an equation for the ellipse whose graph goes through the points (16,0) and (8,6). If your answer is (x^2)/A + (y^2)/B = 1, then A=_____ and B=_____
--------------
Using (16,0) you get: 16^2/A + 0 = 1
A/16^2 = 1
A = 16^2
-------------------
Using (8.6) you get: 8^2/A + 6^2/B = 1
Substitute for A and simplify to get:
--------
(1/2)^2 + 36/B = 1
36/B = 3/4
B/36 = 4/3
B = 48
-----------------
Cheers,
Stan H.